### v2

# Evaluating Sample Utility for Data Selection by Mimicking Model Weights

Abstract
Foundation models are trained on large-scale web-crawled datasets, which often contain noise, biases, and irrelevant information. This motivates the use of data selection techniques, which can be divided into model-free variants—relying on heuristic rules and downstream datasets—and model-based, e.g., using influence functions. The former can be expensive to design and risk introducing unwanted dependencies, while the latter are often computationally prohibitive. Instead, we propose an efficient, model-based approach using the Mimic Score, a new data quality metric that leverages the weights of a reference model to assess the usefulness of individual samples for training a new model. It relies on the alignment between gradients and a target direction induced by the reference model. Using the derived Mimic Scores, we develop Grad-Mimic, a framework that prioritizes samples for learning, creates effective filters, and automates data selection. Empirically, using Mimic Scores to guide training improves data efficiency, results in consistent performance gains across six image datasets, and includes enhancements to CLIP models. Moreover, Mimic Score-based filters improve upon existing filtering methods, e.g., cutting 4.7 million samples to train better CLIP models while offering accurate estimation of training dataset quality.
1 Introduction
Large-scale web-crawled datasets are fundamental to the success of foundation models, e.g., MM1.5 [1], OpenAI CLIP [2], and GPT-4 [3]. These datasets provide vast quantities of information but also carry noise, biases, and irrelevant content. To mitigate these, data selection—ruling out undesirable samples—has emerged as a critical step in the development pipeline [4, 5]. For example, the FineWeb dataset [6], with 15 trillion tokens, is refined through eight carefully designed filtering steps to ensure high model performance.
Existing data selection methods have several limitations. We can divide these into two categories: those that are independent of a model being trained (or using pretrained models) and those that rely on such a model. Model-free approaches, such as those based on handcrafted heuristics or semantic similarity to downstream datasets [7, 8] often require expensive experimentation to establish selection rules and add unwanted dependencies on additional datasets. Furthermore, these methods can fail to provide fine-grained insights into individual samples and overlook sample utility in training, resulting in suboptimal performance. In contrast, model-based techniques, e.g., training specialized filtering networks  [9, 10, 11, 12] or using data influence functions  [13, 14], introduce complexity and are often computationally expensive.
Figure 1: High-value vs. low-value samples identified by the Mimic score: We present randomly selected samples from the top 5% (first row) and bottom 5% (second row) of web-crawled data, ranked by their Mimic scores. Each image is accompanied by its caption and corresponding CLIP score (higher scores suggest better quality). Mimic scores align closely with CLIP scores. High-value samples generally have detailed captions and coherent visual content, while low-value ones carry short captions and misaligned content.
To address these limitations, we introduce Mimic Score, a new data quality metric that assesses sample’s contribution to weight updates using reference model’s weights. This is an efficient model-based technique. We show that by leveraging pretrained model weights (located in a more optimal part of the weight space), we bypass the need for downstream datasets or specialized model training and can use it as our selection guide. Our approach is based on the alignment between each sample’s negative gradient and a direction pointing toward the reference model in the weight space.
We argue that samples that potentially pull the model in undesirable directions—thereby misguiding weight updates—should be considered low-value and ruled out.
Leveraging model weights to evaluate sample utility offers several advantages. First, with the growing availability of well-trained models, adopting model weights has become more accessible than ever. Second, unlike other reference model-based approaches that require loading a second model and performing full inference passes [15, 16], our method scores each sample directly from the weights, improving efficiency.
Building on the mimic score, we develop Grad-Mimic, a data selection framework that operates in two stages:
  • Stage 1: During training, Grad-Mimic prioritizes which samples to learn based on their mimic scores, improving data efficiency while steering the model toward the reference model’s weight space.
  • Stage 2: After training, Grad-Mimic uses mimic scores to evaluate each sample’s utility across training steps. These assessments are then aggregated to create an ensemble filter, which can further refine the dataset.
We validate Grad-Mimic through diverse experiments. First, in a controlled setting with varying levels of label noise, Grad-Mimic accurately identifies mislabeled samples and estimates training dataset quality. During training, it effectively down-weights undesirable samples and improves data efficiency across six image datasets. Next, we test it on large-scale web-crawled multimodal data by training CLIP models from scratch [2] on 10M and 100M samples from DataComp  [17]. Using publicly available pretrained weights as our reference, mimic scores effectively navigate training. The resulting filter performs better than human-designed ones and can complement CLIP score-based filters [18], lifting CLIP model performance with less training data.
We summarize our contributions as follows:
  • New Data Quality Metric: We propose a novel approach to assess sample utility by computing the alignment between gradients and the vector toward the reference model in the weight space.
  • Automated Data Selection Framework: Using the mimic score, we present Grad-Mimic, a framework that effectively identifizes useful samples to improve data efficiency and automates effective data selection.
  • Dataset Quality Assessment: The mimic score, along with its resulting filters, improves existing filtering strategies and provides reliable estimation of dataset quality.
2 Related Work
Figure 2: Grad-Mimic Two-stage Workflow: In Stage 1, Grad-Mimic uses the vector pointing toward the reference model’s weight space to measure alignments with each sample’s negative gradients. Then it adapts alignment scores to reweight gradients that contributed to the weight updates. In Stage 2, derived alignment scores—mimic scores—are used to identify low-value samples and build an ensemble filter, which can be used for noise detection, dataset quality estimation, and training dataset discovery.
Data Selection.
Data selection techniques can be categorized into group- and sample-level approaches. Group-level approaches focus on optimizing data domain mixtures to curate high-quality datasets [19, 20, 21]. Sample-level methods, which are the focus of Grad-Mimic, aim to provide a score measuring sample utility and filter out unhelpful samples. Prior research has explored various gradient-based strategies, such as identifying impactful samples through gradient magnitudes [22], analyzing gradient similarities across batches [23], and selecting key samples that can capture full training [24, 25]. These methods can struggle with large amounts of noise in batch gradients. Grad-Mimic overcome this with the help of a reference model.
Data Curation for Multimodal Models.
Multimodal models rely on large-scale web datasets [26, 17, 27], which are often naturally noisy and require careful curation. Previous approaches have included selecting samples based on semantic similarity to the downstream evaluation datasets [7, 8], using semantic deduplication [28], developing specialized filtering networks [9, 10, 11, 12], and using data influence functions [13, 14]. While effective, they require access to target datasets, introduce training complexities, and can be expensive in computation. Grad-Mimic overcomes these and offers a more efficient alternative: using public available model weights to identify useful samples.
Weak Supervision.
Weak supervision is an emerging paradigm for constructing labeled datasets by combining multiple noisy label estimates [29, 30, 31, 32]. These estimates can come from sources such as heuristic labeling rules, domain knowledge, or pretrained models [33, 34], often encoded as labeling functions. The outputs of these labeling functions are modeled and then aggregated to produce final labels [29]. Weak supervision has demonstrated success across diverse domains [35, 36, 37, 38, 39]. Most prior works focus on weak label aggregation to construct datasets, Grad-Mimic leverages weak supervision for a new purpose: modeling sample utility assessments for a combined filtering decision.
3 Evaluating Sample Utility
We quantify samples by their contributions to the learning process. Our principle is that samples that potentially pull the model in an undesirable direction, thereby misdirecting weight updates, should be considered low-value.
We start with setup and notation, then explain how the scoring metric is derived and used in our data selection framework, Grad-Mimic.
Notation.
Let $D=\{s_{i}\}_{i=1}^{n}$ denote a dataset of $n$ samples drawn from a distribution $\mathcal{D}$ supported on the space $\mathcal{S}$. At training step $t$, model parameters $\theta_{t}$, are iteratively optimized using the dataset $D$. While our framework supports various training settings, we focus on supervised learning for clarity. We assume $\mathcal{S}=\mathcal{X}\times\mathcal{Y}$, where $\mathcal{X}$ is the input space and $\mathcal{Y}$ is the label space. Each sample $s_{i}$ can be expressed as $(x_{i},y_{i})$, where noise may be present either in the instance $x_{i}$ or in the label $y_{i}$. The empirical loss across samples is defined as $\frac{1}{n}\sum_{i=1}^{n}\ell(x_{i},y_{i})$, and each sample’s gradient with respect to the model parameters $\theta_{t}$ is written as: $g_{i,t}:=\nabla_{\theta_{t}}\ell(x_{i},y_{i})$. A standard update to model parameters is $\theta_{t+1}:=\theta_{t}-\eta\sum_{i=1}^{b}g_{i,t}$, where $\eta$ is the learning rate and $b$ is the batch size.
3.1 Mimic Score Calculation
To evaluate whether a sample’s gradient pulls the model in an undesirable direction, we use the weights of a pretrained reference model as our guide. These reference weights, denoted by $\theta_{\text{ref}}$, reside in a more optimal part of the weight space, such that $\ell(\theta_{\text{ref}})<\ell(\theta_{t})$. We use the vector from the current model’s weight space $\theta_{t}$ to $\theta_{\text{ref}}$ to measure each sample utility in approximating a better weight configuration ($\theta_{\text{ref}}$). These weights can be layer-specific, e.g., model weights in the last layer, which usually store more informative features [40, 41].
Reference weights can be obtained in two ways: by training a model on the dataset to achieve a desired performance level, if resources permit, or more efficiently, by using publicly available well-trained models, which are now more accessible and can eliminate the need for direct data access or possible costly training.
The vector pointing toward reference model’s weights, at training step $t$, is represented by $v_{t}:=\theta_{\text{ref}}-\theta_{t}$. We examine how each sample’s negative gradient $-g_{i,t}$, intended for updating model weights, aligns with vector $v_{t}$. We measure this alignment degree by considering both the direction and magnitude of the negative gradient. Specifically, we compute the projection length of $-g_{i,t}$ onto $v_{t}$, yielding an alignment score $m_{i,t}$, computed as follows
[…]
This alignment score, named the mimic score, reflects how much a sample can drive the model closer to the reference model. A sample having a lower mimic score suggests it has limited utility in guiding model updates, making it a potential candidate for exclusion from future training.
4 Grad-Mimic: Data Selection Framework
Building on this simple but effective metric, we propose Grad-Mimic, a two-stage data selection framework that first prioritizes which samples to learn and then identifies useful samples to design an effective filter. The complete Grad-Mimic workflow is illustrated in Fig. 2.
4.1 Stage 1: Reweighted Gradient Updates
We first use mimic scores to aid model training by re-weighting sample gradients. Unlike standard gradient descent, which assigns equal weight to all the samples in the mini-batch, Grad-Mimic uses mimic scores to amplify helpful samples and down-weight unhelpful ones.
To achieve this, each sample’s mimic score is first normalized using the softmax function with a temperature parameter, $\tau$. The normalized score for a sample $s_{i}$ in a batch of size $b$ is computed as
$\overline{m}_{i,t}=\frac{e^{m_{i,t}/\tau}}{\sum_{j=1}^{b}e^{m_{j,t}/\tau}}.$ (2)
Then, the weight update step in Grad-Mimic is modified as
$\theta_{t+1}:=\theta_{t}-\eta\sum_{i=1}^{b}\overline{m}_{i,t}\cdot g_{i,t}.$ (3)
The temperature $\tau$ controls the sensitivity of sample reweighting, allowing us to adjust how sharply the model prioritizes samples. A lower temperature results in a more aggressive focus on learning the most aligned samples, while a higher temperature encourages Grad-Mimic converges to standard gradient descent.
4.2 Stage 2: Automated Filter Design
The second stage of Grad-Mimic uses computed mimic scores to identify valuable samples, creating a combined filter based on sample utility assessments.
4.2.1 Sample Identification
Grad-Mimic first gathers normalized mimic scores for each sample at every training step. These scores indicate the sample contributions over training, which allows us to decide whether to retain or discard a sample. Grad-Mimic supports several sample identification methods:
  • Threshold-based: A sample is chosen if its mimic score exceeds a defined threshold. This threshold could be set as $1/b$ (indicating greater than uniform weight).
  • 1D Clustering: Samples are categorized into two groups using clustering such as $k$-means clustering [42] or Gaussian Mixture Models (GMMs), allowing more automated identification based on their assigned cluster.
  • Top-k Percent: Samples are ranked by their mimic score, and the top-$k$ percent are chosen. The value of $k$ can be adjusted based on the available training budget.
After identification, normalized mimic scores are binarized, and samples are categorized into “retain” or “discard” at each training step. We treat these assigned groups as each training step’s votes and then combine for a final filter.
4.2.2 Filter Output Aggregation
Although using the assessment at a training step can refine the dataset, relying on such isolated filters risks overlooking training dynamics. Moreover, evaluating a sample’s utility across all training stages may yield complementary signals.
To address this, Grad-Mimic uses weak supervision techniques [29, 30, 31, 32], which are usually used for constructing datasets from noisy signals, to combine filter outputs across training steps. We start by learning a model to evaluate the reliability of each step’s assessment [31]. Once established, this model aggregates filtering decisions into a high-quality ensemble filter. Through aggregation, Grad-Mimic mitigates assessment noise and captures training dynamics. We employ the Snorkel framework [30], a widely used method in the weak supervision community.
Ultimately, we can use this ensemble filter to select high-value samples for a refined dataset. Moreover, we can use it to estimate training dataset quality by either analyzing the proportion of retained samples or by averaging samples’ mimic scores as an overall metric. We summarize the steps of Grad-Mimic in Algorithm 1 and detailed in Appendix A.
5 Experiments
We evaluate the effectiveness of Grad-Mimic using two experimental setups across datasets of varying scales and domains. We first test in a controllable setting, adding noise into sample labels (Sec. 5.1), followed by evaluations on large-scale web-crawled datasets (Sec. 5.2). Our goals are to validate the following claims in both setups:
  • Enhanced Data Efficiency: A pretrained reference model serves as a reliable guide, enabling mimic scores to prioritize samples that improve model performance.
  • Accurate Sample Identification: Grad-Mimic can detect noisy samples, construct effective filters, and automate data selection.
  • Dataset Quality Assessment: Mimic scores and their resulting filters provide reliable estimates of dataset quality.
Full Fine-tuning DTD Flowers102 STL10 Oxford-IIIT Pet CIFAR10 CIFAR100 Average
Ref. Model 65.59 91.01 93.20 82.88 94.89 81.34 84.82
Noise Level 0.4 0.5 0.6 0.4 0.5 0.6 0.4 0.5 0.6 0.4 0.5 0.6 0.4 0.5 0.6 0.4 0.5 0.6 0.4 0.5 0.6
Mini-batch SGD 46.91 42.02 33.24 65.54 55.90 39.84 75.35 58.57 49.98 70.43 64.62 55.63 90.31 86.69 82.01 69.42 63.77 59.28 69.66 61.93 53.33
GraNd 29.73 8.83 15.43 10.80 2.78 27.92 78.39 52.32 64.30 56.83 45.38 45.43 87.56 83.79 80.04 18.65 8.83 7.84 46.99 33.66 40.16
AGRA 46.28 37.29 32.39 0.62 0.93 0.82 84.35 79.94 78.17 76.97 77.11 69.39 89.23 86.79 68.09 11.96 8.44 6.88 51.57 48.42 42.62
Grad-Match 48.88 42.66 33.24 37.92 54.94 42.62 83.86 81.16 77.18 74.35 61.73 51.16 44.48 35.69 37.59 69.41 66.21 20.09 58.82 57.07 43.65
Grad-Mimic 49.20 42.82 33.83 68.58 56.46 44.27 72.06 71.85 83.09 81.98 78.30 73.34 90.52 89.07 66.41 73.97 74.31 24.02 72.72 68.60 54.16
Linear Probing DTD Flowers102 STL10 Oxford-IIIT Pet CIFAR10 CIFAR100 Average
Ref. Model 60.00 56.98 97.41 89.45 94.18 77.89 79.32
Noise Level 0.4 0.5 0.6 0.4 0.5 0.6 0.4 0.5 0.6 0.4 0.5 0.6 0.4 0.5 0.6 0.4 0.5 0.6 0.4 0.5 0.6
Mini-batch SGD 51.91 47.71 44.44 28.64 21.50 15.43 96.26 95.49 93.56 87.33 85.55 83.51 92.86 92.14 91.19 75.56 74.38 73.25 72.09 69.46 66.40
GraNd 36.22 30.59 25.32 18.23 13.90 10.49 68.70 59.55 49.68 69.47 60.02 48.13 88.02 85.61 81.48 71.60 70.51 69.27 58.71 53.36 47.40
AGRA 41.81 36.28 31.49 41.81 36.28 31.49 96.19 95.15 93.11 85.25 82.53 78.39 92.51 92.01 90.99 72.50 71.30 69.69 71.68 68.93 65.86
Grad-Match 51.81 47.55 41.33 27.00 20.21 14.90 96.06 95.34 93.44 86.86 85.75 83.18 92.62 92.04 91.17 75.39 74.46 73.18 71.62 69.23 66.20
Grad-Mimic 54.68 54.10 50.43 42.75 37.10 31.71 97.16 97.00 96.90 88.80 88.25 87.14 94.15 93.92 93.80 77.24 76.82 76.05 75.80 74.53 73.01
Table 1: Stage 1 Results in Simulation Experiment: Using mimic scores can effectively down-weight mislabeled samples during training, improving data efficiency and closing the performance gap to reference model.
DTD Flowers102 STL10 Oxford-IIIT Pet CIFAR10 CIFAR100
Noise Level 0.4 0.5 0.6 0.4 0.5 0.6 0.4 0.5 0.6 0.4 0.5 0.6 0.4 0.5 0.6 0.4 0.5 0.6
Threshold ($1/32$) 85.13 88.91 91.82 90.93 94.9 96.92 90.31 94.11 97.21 94.41 96.86 97.93 98.37 98.19 97.19 97.66 97.57 95.80
1D $k$-means 77.13 76.68 77.17 84.45 87.62 87.47 78.90 83.04 78.82 95.80 95.49 93.73 98.62 98.17 97.56 98.10 97.46 96.62
GMM 97.85 96.72 95.68 98.39 96.96 94.21 97.96 97.16 95.69 98.04 97.07 95.86 95.70 92.89 88.62 95.86 94.31 92.55
Table 2: Stage 2 Results in Simulation Experiment: We report detection results using F1-score metric. Grad-Mimic consistently achieves precise detection of mislabeled samples across datasets and proposed identification methods. Notably, its performance remains high and robust regardless of the noise level.
5.1 Simulating Mislabeled Samples
Setups.
We begin with a controlled experiment by adding various levels of label noise to six image classification datasets. They are DTD [43], Flowers102 [44], STL10 [45], OxfordIIIT Pet [46], CIFAR10, and CIFAR100 [47]. We fine-tune the ViT-B/16 model [48] on these noisy datasets under two configurations: linear probing, where only the final layer is tuned, and full fine-tuning, where gradients of all model parameters are reweighted based on mimic scores to maximize the impact of useful samples. We normalize mimic scores with a temperature of 0.5 and use a batch size of 32. We simulate pretrained reference models by training ViT-B/16 models on the noise-free version of each dataset and use the last-layer weights as our reference to navigate training on the noisy datasets. We detail more training configurations in Appendix B.1.
After training, we identify samples in two ways: setting one over batch size ($1/b$) as our threshold and using $k$-means and GMM to cluster. Then we aggregate filter outputs across training steps using Snorkel framework [30].
Figure 3: Mimic Score Distribution: Extracted from CIFAR100 dataset with a noise level of 0.5. The distributions clearly separate, aligning with the correctness of the labels.
Expected Results.
We aim to validate that a pretrained reference model’s weights can serve as a reliable guide. Moreover, we expect that mimic scores can help identify low-value, mislabeled samples.
Baselines.
We compare Grad-Mimic’s training method (Stage 1) with other approaches that use gradient information for sample prioritization. We consider: (i) Mini-batch SGD, the standard method, updates weights by averaging gradients within the mini-batch. (ii) GraNd [22] reweights samples based on their gradient norm, prioritizing data that induce greater changes. (iii) AGRA [23] computes the cosine similarity of each sample gradient to an average gradient of another random batch, then excludes outlier samples. (iv) Grad-Match [24] adapts gradients by solving a subset selection problem to identify key samples within each batch. Furthermore, we compare Grad-Mimic with Rho-Loss [15], a reference model-based approach that prioritizes training on worth learning samples, as well as influence function-based methods. Results are presented in Appendix C. Grad-Mimic outperforms both techniques while additionally offering efficiency advantages.
Stage 1 Results.
We present Grad-Mimic’s two-stage results separately. Stage 1 results, shown in Table 1, report testing accuracy across various methods and noisy datasets using both training configurations. These results demonstrate that Grad-Mimic effectively identifies and down-weights mislabeled samples during training, enhancing denoising and narrowing the performance gap with the reference model. Notably, in the full fine-tuning setup—where we only target the last layer to mimic—Grad-Mimic successfully guides weight updates for the entire model, ultimately outperforming other methods. This highlights its computational efficiency, as it avoids requiring sample gradients from the full set of model weights. Additionally, we explore the impact of reference model quality in Appendix C, further strengthening our approach. These findings validate the idea that using the direction induced by a better-performing model can guide model training and improve data efficiency.
Stage 2 Results.
Next, we demonstrate Grad-Mimic’s capability to detect mislabeled samples. We use F1-scores to report the detection results in Table 2. Grad-Mimic accurately identifies mislabeled data across datasets and remains robust when facing different noise levels.
We use CIFAR100 dataset as an example and visualize computed mimic score distribution at each training step in Figure 3. In this dataset, half of the labels are flipped. We can see their distributions clearly separate into two groups, confirming that mimic scores are informative signals to identify samples. Despite variations over training, these dynamics can be captured in our aggregation step.
We use the ensemble filter from each dataset created by GMM clustering to estimate training dataset quality using the proportion of retained data. The results, displayed in Figure 4, highly correspond to the presence of label noise, with a Pearson correlation of -0.903. These demonstrate mimic score’s effectiveness for dataset quality estimation.
5.2 Data Selection in Large-scale Web Datasets
Setups.
Next, we test Grad-Mimic in a more challenging setting using million-scale web-crawled datasets. We use the small-scale and medium-scale DataComp datasets [17], which contain approximately 10 million and 100 million image-caption pairs [^1] We identified that some URLs provided by DataComp dataset are now broken. This means that our results might not be comparable to previous approaches on the DataComp Leaderboard. See here for details., respectively, to train CLIP model from scratch [2]. In these datasets, noise is naturally presented in the data. We follow the training setup from DataComp [49, 17] and use publicly available pretrained CLIP model’s weights as our reference  [^2] https://huggingface.co/laion/CLIP-ViT-B-32-DataComp.XL-s13B-b90K. This reference model, trained on the DataComp-1B dataset (1.4 billion samples), represents the best-performing model accessible to us. It serves as a proxy for the ideal reference point for scenarios where training such powerful models is infeasible. We use the final MLP layer in the text and image encoders respectively as our target to mimic. We evaluate the performance of resulting model using DataComp benchmark, which includes 38 diverse image classification and retrieval tasks. Other details can be found in Appendix B.2.
In Stage 2, we use top-$k$ percent identification to create the final filter, subsequently training a second CLIP model on the curated dataset to validate resulting filter’s effectiveness.
Scale Training Method Mimic Layer $\theta_{\text{ref}}$ Temperature $\tau$ ImageNet
ImageNet
dis. shifts
VTAB Retrieval
Average over
38 datasets ($\uparrow$)
Vanilla Training 0.026 0.035 0.139 0.114 0.131
0.05 0.026 0.035 0.147 0.114 0.135
Last MLP Layer in Text Encoder 0.5 0.027 0.035 0.152 0.114 0.139
0.05 0.026 0.037 0.162 0.118 0.146
Small Grad-Mimic Last MLP Layer in Image Encoder 0.5 0.027 0.035 0.160 0.117 0.145
Vanilla Training 0.171 0.148 0.253 0.217 0.254
Medium Grad-Mimic
Last MLP Layer
in Image Encoder
0.05 0.169 0.151 0.262 0.216 0.258
Table 3: Stage 1 Results in DataComp Experiment: On both dataset scales, with the aid of publicly available pretrained weights, Grad-Mimic yields higher CLIP model performance. Full table with various temperature settings is placed in Table 6 in Appendix  C.
Expected Results.
We anticipate mimic scores help large-scale training focus on high-value samples and can serve as instrumental metric for automating effective data curation.
Baselines.
In the first stage, we compare Grad-Mimic to vanilla training, where each sample contributes in the weight update equally. After training, we compare our mimic score-based filter against the following methods: (1) No Filtering: using the entire training pool, (2) Basic Filtering [17]: selecting samples based on predefined criteria like caption length, image size, and caption language—representing a human-designed filter, (3) CLIP Score [18]: selecting top-$k$ percent samples based on embedding alignment between images and captions. Scores are computed by OpenAI’s CLIP ViT-B/32 model pretrained on over 400 million samples [2]. We see CLIP score as a strong baseline, as it is prebuilt through heavy training using hundreds of millions of samples.
Stage 1 Results.
Table 3 presents pretraining results for CLIP models on both dataset scales. Grad-Mimic consistently outperforms standard training across all temperature settings. Full results using different temperatures are provided in Appendix C. Interestingly, we find that mimicking the last layer of the image encoder yields more performance gains compared to targeting the text encoder. These results highlight Grad-Mimic’s success in using online pretrained weights as a reference to prioritize high-value samples for learning.
Scale Filtering Strategy Dataset Size ImageNet
ImageNet
dis. shifts
VTAB Retrieval
Average over
38 datasets ($\uparrow$)
No Filtering 10.7M 0.026 0.035 0.139 0.114 0.131
Basic Filtering 3M 0.031 0.040 0.159 0.115 0.141
Mimic Score Top 30% 3M 0.028 0.037 0.151 0.107 0.135
Mimic Score Top 35% 3.8M 0.026 0.037 0.158 0.107 0.143
Mimic Score Top 40% 4.2M 0.028 0.035 0.151 0.110 0.139
CLIP Score (B/32 Top 30%) 3M 0.051 0.054 0.183 0.118 0.165
CLIP Score $\setminus$ Mimic Score Bottom 15% 2.7M (0.3M $\downarrow$) 0.047 0.053 0.174 0.106 0.163
Small CLIP Score $\setminus$ Mimic Score Bottom 20% 2.6M (0.4M $\downarrow$) 0.047 0.052 0.173 0.113 0.164
No Filtering 101.9M 0.026 0.035 0.139 0.114 0.254
Basic Filtering 30M 0.209 0.176 0.264 0.232 0.267
Mimic Score Top 30% 29M (1M $\downarrow$) 0.192 0.163 0.269 0.218 0.268
CLIP Score (B/32 Top 30%) 30M 0.275 0.229 0.332 0.246 0.321
CLIP Score $\setminus$ Mimic Score Bottom 5% 28.1M (1.9M $\downarrow$) 0.270 0.227 0.339 0.246 0.323
CLIP Score $\setminus$ Mimic Score Bottom 10% 27.7M (2.3M $\downarrow$) 0.274 0.226 0.334 0.243 0.323
Medium CLIP Score $\setminus$ Mimic Score Bottom 30% 25.3M (4.7M $\downarrow$) 0.264 0.223 0.330 0.244 0.322
Table 4: Stage 2 Results in DataComp Experiment: Mimic score-based filters perform better than basic filtering and complement CLIP score-based filters by removing low-value samples. Symbol $\setminus$” denotes the exclusion of curated datasets.
((a))
((b))
Figure 4: Mimic scores offer precise dataset quality estimation: Left figure computes the portion of retained samples to evaluate dataset quality in simulation experiments. Right figure takes the average on mimic scores in each filtered dataset for estimation. [S/M] denotes the scale of DataComp dataset—small [S] or medium [M]—that the filter is designed on.
Stage 2 Results.
We use the derived mimic scores from our best-performing model in Stage 1 to design filters and evaluate their effectiveness. Results are presented in Table 4. Grad-Mimic automates the selection, outperforming basic filtering strategy at both scales. Additionally, we remove low-value samples identified by mimic score to improve CLIP score-based filter (B/32 Top 30%). This complementary approach not only enhances model performance but also improves data efficiency by reducing the training set size, specifically cutting down 4.7 million samples in the medium-scale dataset.
Lastly, we gather mimic scores from datasets that are curated by various CLIP score-based filters. We include the top-ranked approach during our time of experimentation [50], which uses an orthogonal approach based on object detection models. We average samples’ mimic scores in each curated dataset and compare with their corresponding performance gains. The results, shown in Figure 4, reveal a positive alignment with a Pearson correlation of 0.958, demonstrating mimic score as a reliable metric for dataset quality estimation.
Sample Analysis.
We analyze high-value and low-value samples ranked by mimic score. In Figure 5, we compare token usage in captions. We find that bottom-ranked samples often carry captions that are either too short (1-2 words) or excessively long (over the maximum token limit), unlike top-ranked samples. Moreover, mimic score-based filter effectively identifies low-value samples, such as Captcha-like images in the small-scale dataset and empty or incoherent images in the medium-scale dataset (see examples in Figure 1). These filtering patterns align with human intuition but are automatically captured by Grad-Mimic.
We visualize the top and bottom 30% of samples along with their CLIP scores in Figure 6. High-value samples align with higher CLIP scores, while low-value ones are more concentrated in the lower range (half below 0.2). Furthermore, we calculate their correlation on random samples. We normalize both scores into the range 0 to 1 for clarity The observed positive correlation supports mimic scores in identifying high-quality samples.
Figure 5: The filtering logic behind Grad-Mimic aligns with handcrafted heuristics, particularly in caption length. Low-value samples (left figure) tend to have captions that are either too short or exceed the maximum token length, i.e., 77.
((a))
((b))
Figure 6: Mimic score distribution and correlation to CLIP score.
Filtering Strategy Random Selection Mimic Score Selection
Jaccard
Similarity
Percentage of
Overlap
Jaccard
Similarity
Percentage of
Overlap
CLIP-ViT B/32 Top 30% 0.149 0.258 0.232 0.376
CLIP-ViT L/14 Top 30% 0.150 0.260 0.236 0.382
DataComp’23 Top-ranked 0.166 0.284 0.271 0.426
Table 5: Mimic score can help in predicting whether a sample was used to train the reference model.
Predicting Pretraining Dataset
We explore an interesting application of mimic score: predicting the pretraining dataset. Specifically, we ask: how accurately can Grad-Mimic check whether a given sample was used to train a reference model? We test this hypothesis on the small-scale DataComp dataset. We first apply various pre-built filters to curate datasets and train CLIP models on each. Then, we target their final-layer weights to mimic and apply Grad-Mimic to train on the entire data pool. Our goal is to see whether the top-ranked samples (matched in size to the curated datasets) identified by mimic score appear in their training dataset.
We evaluate our approach against random selection using Jaccard similarity and percentage of overlap. As shown in Table 5, mimic score-based selection identifies more samples used to train the reference model compared to random sampling. We achieve a 42.6% overlap with DataComp’23 best-performing filtered dataset by simply mimicking final-layer weights without direct access to their filtering steps.
6 Conclusion
We introduce Mimic Score, an efficient data quality metric that assesses samples’ utility in model training. We quantify samples by measuring the alignment between gradients and the vector toward pretrained reference weights. Using this, we develop Grad-Mimic to automate data selection. Empirically, Grad-Mimic effectively prioritizes samples for learning, improving data efficiency and model performance. Moreover, Mimic Score aids in creating effective filters and estimating dataset quality. Resulting filters outperform human-designed ones and exhibit a strong correlation with dataset quality and performance gains.
7 Impact Statements
Efficient data selection is increasingly important for training foundation models. We demonstrate Grad-Mimic’s effectiveness in identifying and removing low-value samples, which enhances data efficiency and resulting model performance. We do not foresee explicit negative social impacts arising from Grad-Mimic. However, if the reference model used to evaluate sample utility is suboptimal, there remains a risk of refining low-quality datasets (we have an analysis of the impact of reference model quality in Appendix C). We believe that careful selection and validation of this reference model, a key area of future research, will minimize this risk and enable the creation of improved datasets.
8 Acknowledgments
We thank Kwei-Herng Lai, Prasad Gabbur, Meng Cao, Ryan Jia, Russ Webb, Vedant Joshi, Albert Ge, John Cooper, Harit Vishwakarma, Dyah Adila, Changho Shin, Satya Sai Srinath Namburi, Zhiqi Gao, Avi Trost, and Gabe Orlanski for their helpful feedback and valuable discussion.
Appendix A Grad-Mimic Algorithm
We summarize the algorithm steps in Grad-Mimic for computing mimic scores and updating model weights. Our algorithm supports various training paradigms, including both supervised learning and self-supervised learning. We validate its effectiveness through controlled simulation experiments (Sec. 5.1) and large-scale training of CLIP models on web-crawled datasets (Sec. 5.2). Furthermore, our experiments encompass a range of training configurations, such as full-parameter fine-tuning, linear probing, and training models from scratch.
Algorithm 1 Grad-Mimic (Stage 1)
reference model $\theta_{\text{ref}}$, new model $\theta$, dataset $D$, number of training steps $T$, batch size $b$, learning rate $\eta$, temperature $\tau$
normalized mimic scores $\overline{m}_{i,t}$, a well-trained model $\theta_{T}$
Initialize new weights $\theta_{0}$
for $t\in[T]$ do
     uniformly sample a batch $\{s_{i}\}_{i=1}^{b}$ from $D$
     $v_{t}\leftarrow\theta_{\text{ref}}-\theta_{t}$triangleright compute the vector as a guide
     for $i\in[b]$ do
         $g_{i,t}\leftarrow\nabla_{\theta_{t}}\ell(s_{i})$triangleright compute per-sample gradient
         $m_{i,t}\leftarrow\frac{\langle-g_{i,t},v_{t}\rangle}{\|v_{t}\|}$triangleright compute mimic score
     end for
     $\overline{m}_{:,t}=\frac{e^{m_{:,t}/\tau}}{\sum_{j=1}^{b}e^{m_{j,t}/\tau}}$triangleright normalize mimic scores
     $\theta_{t+1}:=\theta_{t}-\eta\sum_{i=1}^{b}\overline{m}_{i,t}\cdot g_{i,t}$triangleright reweight gradients
end for
Appendix B Experimental Details
We provide more details about our training setups and computational resources.
B.1 Simulation Experiments
In the experiments of simulating mislabeled samples (Sec. 5.1), we fine-tune ViT-B/16 models pretrained on the ImageNet-21k dataset [51]. Each dataset is trained with a batch size of 32, a learning rate of 1e-4, the AdamW optimizer, and for 5 epochs. We evaluate Grad-Mimic using various temperature values $\tau$ (1.0, 0.9, 0.8, 0.7, 0.6, and 0.5) and report the performance at $\tau=0.5$ in the main paper. Results for other temperature values are presented in Table 7. To simulate the reference model, we train ViT-B/16 models on noise-free versions of the datasets but with different random seeds for weight initialization. These experiments were performed on a Nvidia Tesla A100.
B.2 DataComp Experiments
We adopt the training setup from DataComp [49, 17] and train CLIP models from scratch using contrastive objective over image-caption pairs. Each CLIP model is trained for 5 epochs with a batch size of 4096. The total number of seen samples is 12.8M for the small-scale dataset and 128M for the medium-scale dataset. These experiments run on 8 Nvidia Tesla A100s.
We test Grad-Mimic using different parts of weights in the reference model, specifically the final MLP layer in the image and text encoders. Grad-Mimic’s performance is evaluated using 38 diverse downstream tasks [17] with various temperature values $\tau$ (0.03, 0.05, 0.07, 0.3, and 0.5). We report results at $\tau=0.5$ and $\tau=0.05$ in the main body in Table 3, but full results are presented in Table 6.
Scale Training Method Mimic Layer $\theta_{\text{ref}}$ Temperature $\tau$ ImageNet
ImageNet
dist. shifts
VTAB Retrieval
Average over
38 datasets ($\uparrow$)
Vanilla Training 0.026 0.035 0.139 0.114 0.131
0.03 0.026 0.035 0.153 0.115 0.136
0.05 0.026 0.035 0.147 0.114 0.135
0.07 0.026 0.035 0.151 0.116 0.135
0.3 0.026 0.034 0.154 0.112 0.137
Last MLP Layer in Text Encoder 0.5 0.027 0.035 0.152 0.114 0.139
0.03 0.025 0.035 0.149 0.114 0.133
0.05 0.026 0.037 0.162 0.118 0.146
0.07 0.027 0.036 0.166 0.118 0.145
0.3 0.026 0.035 0.161 0.114 0.144
Small Grad-Mimic Last MLP Layer in Image Encoder 0.5 0.027 0.035 0.160 0.117 0.145
Vanilla Training 0.171 0.148 0.253 0.217 0.254
Medium Grad-Mimic
Last MLP Layer
in Image Encoder
0.05 0.169 0.151 0.262 0.216 0.258
Table 6: Full Stage 1 Results in DataComp Experiment: On both dataset scales, with the aid of publicly available pretrained weights, Grad-Mimic consistently improves CLIP model performance across all temperature settings.
Appendix C Ablation Studies
Temperature $\tau$ DTD Flower102 Oxford-IIIT Pet CIFAR10 CIFAR100
Mini-Batch SGD 52.8 35.6 87.9 93.2 76.2
Grad-Match 53.1 32.1 87.4 93.0 76.1
1.0 56.5 44.9 89.0 94.1 77.0
0.9 56.4 45.3 89.1 94.1 77.1
0.8 56.3 45.8 89.1 94.1 77.2
0.7 56.3 46.1 89.2 94.2 77.2
Grad-Mimic 0.6 56.0 46.4 89.3 94.2 77.3
Table 7: Stage 1 Results in Simulation Experiment using Various Temperature: Lower temperature makes the model focus more on training with high-value samples, resulting in a higher testing accuracy.
The Choice of Temperature.
We test Grad-Mimic using different temperature values and compare them to baseline methods (Mini-batch SGD and Grad-Match [24]). The results in the simulation experiments are presented in Table 7. In these datasets, we set the noise level to 0.3 and fine-tune ViT-B/16 model under linear probing configuration. Grad-Mimic outperforms the baseline methods in all temperature settings. Lower temperatures yield better testing accuracy, as they normalize mimic scores in the way that encourages the model to focus more on high-value samples during training.
DTD Flowers102 STL10 Oxford-IIIT Pet CIFAR10 CIFAR100 Average
Ref. Model 60.00 56.98 97.41 89.45 94.18 77.89 79.32
Noise Level 0.4 0.5 0.6 0.4 0.5 0.6 0.4 0.5 0.6 0.4 0.5 0.6 0.4 0.5 0.6 0.4 0.5 0.6 0.4 0.5 0.6
Mini-Batch SGD 51.91 47.71 44.44 28.64 21.50 15.43 96.26 95.49 93.56 87.33 85.55 83.51 92.86 92.14 91.19 75.56 74.38 73.25 72.09 69.46 66.40
Rho-Loss 54.10 52.39 48.46 41.55 35.71 30.02 97.10 96.83 96.49 88.53 87.79 86.86 94.07 93.83 93.82 76.82 76.05 75.93 75.36 73.77 71.93
Grad-Mimic 54.68 54.10 50.43 42.75 37.10 31.71 97.16 97.00 96.90 88.80 88.25 87.14 94.15 93.92 93.80 77.24 76.82 76.05 75.80 74.53 73.01
Table 8: Grad-Mimic outperforms other reference model-based methods: Using the last-layer weights of reference model is more effective to prioritize samples for learning while reducing memory usage and computational overhead.
Comparison to Reference Model-Based Approaches.
We compare Grad-Mimic to Rho-Loss [15], a method that prioritizes samples for learning based on excess loss, computed as the difference between the loss on the current model $\ell(\theta_{t};s_{i})$ and the loss on the pretrained reference model $\ell(\theta_{\text{ref}};s_{i})$. We follow our simulation experiment setup and evaluate the performance of fine-tuned models on the testing datasets.
As shown in Table 8, Grad-Mimic consistently outperforms Rho-Loss across all datasets and noise levels. Furthermore, Grad-Mimic offers efficiency advantages. Unlike Rho-Loss, which requires loading the entire reference model and performing additional inference passes, Grad-Mimic computes mimic scores directly from the reference model’s weights. In our experiments, we demonstrate that even using only a subset of the weights (e.g., the last-layer weights) is sufficient, further reducing memory usage and improving computational efficiency.
Figure 7: Grad-Mimic can outperform influence function-based methods: Using model weights as a guide offer advantages on their effectiveness and greater accessibility.
Comparison to Influence Function-based Approaches.
Data influence functions provide a principled way to evaluating the impact of individual training samples on model parameters [52, 53], but their computational cost often limits practical use. Several studies have explored efficient approximations that estimate sample influence by measuring gradient alignment with a validation dataset [54, 55]. We adapt these methods within Grad-Mimic framework by replacing the vector induced by reference model’s weights ($v_{t}$) with gradients computed from validation dataset. We evaluate these adaptations on six image datasets by adding 50% label noise and assess the impact of validation set size by varying the number of samples drawn from each class (1, 5, 10, 15, and 20).
Results in Figure 7 show that while larger validation sets improve data efficiency by enabling model to focus on the influential samples, Grad-Mimic—using pretrained reference weights and their induced direction—consistently outperforms validation dataset-based influence approximations. Moreover, relying on a labeled validation set risk introducing an unwanted dependency. For example, in the out-of-domain Flowers102 dataset, achieving satisfactory performance required over 2000 correctly labeled validation samples—yet still fell short of Grad-Mimic’s results.
Figure 8: Grad-Mimic can be robust to the quality of reference model.
Impact of Reference Model Quality.
Finally, we investigate how the quality of the reference model impacts the effectiveness of Grad-Mimic, particularly in scenarios where only a weaker or distilled model is available. To simulate such scenarios, we inject zero-mean Gaussian noise $(0,\sigma^{2})$ into the reference model’s weights. By varying the variance of the added noise, we analyze the performance of models trained with Grad-Mimic against standard training methods that do not leverage a reference model.
We present this analysis in Figure 8. As expected, increasing noise variance degrades the reference model’s performance, which in turn lowers the effectiveness of Grad-Mimic. However, we observe that even under significant noise, models trained with Grad-Mimic still outperform Mini-batch SGD and can even surpass the performance of the noisy reference model itself. This highlights that while a high-quality reference model is ideal, even a weaker proxy can provide valuable guidance and improve training outcomes over traditional baselines.